# HyPA-RAG: A Hybrid Parameter Adaptive Retrieval-Augmented Generation System for AI Legal and Policy Applications

### v1

@@ 9 Overall Results and Discussion
Our analysis shows that adaptive methods generally outperform fixed baselines, particularly in improving faithfulness and answer quality. Incorporating adaptive parameters such as query rewrites and reranking enhances the system’s ability to provide accurate and relevant responses. While adding a reranker improves correctness, it can slightly reduce the overall correctness score, suggesting a trade-off between precision and answer quality.
The introduction of a knowledge graph maintains correctness but can add complexity, potentially lowering overall response quality. However, combining adaptive parameters with a reranker proves effective in maximizing the proportion of correct responses, even if it doesn’t lead to the highest scores in all metrics.
Overall, these findings highlight the importance of adaptivity and careful parameter tuning to balance different performance aspects, enhancing the system’s capability to handle varied and complex queries effectively.
10 Limitations and Future Work
Figure 4: RAG System Optimisation Feedback Loop
Integrating human feedback into the evaluation loop (see Figure 4) could better align metrics with user preferences and validate performance metrics in real-world settings. Future work should also consider fine-tuning the LLM using techniques like RLHF (Bai et al., 2022), RLAIF (Lee et al., 2023), or other preference optimisation methods (Song et al., 2023). Further, refining the query rewriter (Ma et al., 2023; Mao et al., 2024) and exploring iterative answer refinement (Asai et al., 2023) could enhance metrics like relevancy and correctness.
11 Ethical and Societal implications
@@
12 Acknowledgments
We would like to express our sincere gratitude to Holistic AI for their invaluable research and financial support, which made this project possible. Our deepest appreciation also goes to the Center for Artificial Intelligence at University College London for their continuous encouragement and assistance. Without their collaboration and resources, this work would not have been achievable.
@@ Appendix A Appendix
A.1 HyPA RAG Demo UI
Figure 5: HyPA-RAG Configuration Demo
@@ A.9.3 Concise Prompt
A.11 2-Class Ablation Results

### v2

@@ 9 Overall Results and Discussion
Our analysis demonstrates that adaptive methods outperform fixed baselines, particularly in faithfulness and answer quality. Adaptive parameters, such as query rewrites and reranking, enhance response accuracy and relevance, though reranking may slightly reduce overall correctness scores, indicating a trade-off between precision and quality. Adding a knowledge graph maintains correctness but introduces complexity, potentially lowering response quality. However, combining adaptive parameters with reranking maximizes correct responses, even if it doesn’t achieve the highest scores across all metrics. These findings demonstrate the effect of adaptivity and parameter tuning to balance performance, enabling effective handling of diverse and complex queries. This suggests our system could also apply to other domains where queries demand complex, multi-step reasoning and non-obvious concept relationships. Limitations and future work are detailed in Appendix A.13.
10 Ethical Considerations
@@ Appendix A Appendix
A.1 RAG Demonstration User Interface
(a) Demo Screenshot: Entering the user query and generating a response.
(b) Demo Screenshot: The generated response.
(c) Demo Screenshot: Information on retrieved node metadata and content.
Figure 4: Demo screenshots showing each key stage of the user experience.
@@ A.9.3 Concise Prompt
A.11 3-Class Ablation Results
Method Faithfulness Answer
Relevancy
Absolute
Correctness (1-5)
Correctness
(Threshold=4.0)
$k$ 0.7723 0.7940 4.0409 0.7621
$k$, $Q$ 0.8971 0.7778 4.2528 0.8141
$k$, $Q$ + reranker 0.9098 0.7902 4.2342 0.8178
$k$, $K^{*}$, $S^{*}$ 0.8733 0.7635 4.1227 0.8141
$k$, $K$, $S$ 0.8660 0.7780 4.1822 0.8030
$k$, $K$, $S$ + reranker 0.8821 0.7872 4.1858 0.8178
$k$, $K$, $S$, $Q$ 0.8465 0.7734 4.1338 0.7918
$k$, $K$, $S$, $Q$ + reranker 0.8689 0.7853 4.1859 0.8402
A.12 2-Class Ablation Results
@@
A.13 Future Work and Limitations
Integrating human feedback into the evaluation loop could better align metrics with user preferences and validate performance metrics in real-world settings. Future work should also consider fine-tuning the LLM using techniques like RLHF (Bai et al., 2022), RLAIF (Lee et al., 2023), or other preference optimisation methods (Song et al., 2023). Further, refining the query rewriter (Ma et al., 2023; Mao et al., 2024) and exploring iterative answer refinement (Asai et al., 2023) could enhance metrics like relevancy and correctness.